# Classical and Quantum KMS States on Spin Lattice Systems

### v1

@@ 3.2 Semi-classical limit of quantum KMS condition
$-i\beta\omega_{\infty}^{\beta,\Gamma}\big{(}a_{\Lambda}^{*}\delta_{\infty}(a_{\Lambda})\big{)}\geq-\lim_{j\to\infty}\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}(2j+1)\log\bigg{(}1+\frac{\omega_{j}^{\beta,\Gamma}\big{(}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}(a_{\Lambda}^{*})]\big{)}}{\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}}\bigg{)}\\ =-\lim_{j\to\infty}\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}(2j+1)\log\bigg{(}1+\frac{i}{2j+1}\frac{\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(\{a_{\Lambda},a_{\Lambda}^{*}\}_{B_{\infty}^{\Lambda}})\big{)}}{\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}}\bigg{)}\\ =-i\lim_{j\to\infty}\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(\{a_{\Lambda},a_{\Lambda}^{*}\}_{B_{\infty}^{\Lambda}})\big{)}=-i\omega_{\infty}^{\beta,\Gamma}\Big{(}\{a_{\Lambda},a_{\Lambda}^{*}\}_{B_{\infty}^{\Gamma}}\Big{)}\,.$

### v2

@@ 3.2 Semi-classical limit of quantum KMS condition
$-i\beta\omega_{\infty}^{\beta,\Gamma}\big{(}a_{\Lambda}^{*}\delta_{\infty}(a_{\Lambda})\big{)}\geq-\lim_{j\to\infty}\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}h_{j}^{-1}\log\bigg{(}1+\frac{\omega_{j}^{\beta,\Gamma}\big{(}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}(a_{\Lambda}^{*})]\big{)}}{\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}}\bigg{)}\\ =-\lim_{j\to\infty}\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}h_{j}^{-1}\log\bigg{(}1+\frac{i}{2j+1}\frac{\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(\{a_{\Lambda},a_{\Lambda}^{*}\}_{B_{\infty}^{\Lambda}})\big{)}}{\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(a_{\Lambda}^{*})Q_{j}^{\Lambda}(a_{\Lambda})\big{)}}\bigg{)}\\ =-i\lim_{j\to\infty}\omega_{j}^{\beta,\Gamma}\big{(}Q_{j}^{\Lambda}(\{a_{\Lambda},a_{\Lambda}^{*}\}_{B_{\infty}^{\Lambda}})\big{)}=-i\omega_{\infty}^{\beta,\Gamma}\Big{(}\{a_{\Lambda},a_{\Lambda}^{*}\}_{B_{\infty}^{\Gamma}}\Big{)}\,.$